# On the Fundamental Impossibility of Hallucination Control in Large Language Models

### v5

@@ 4 Theoretical Framework
This section develops the mathematical foundations necessary for proving the impossibility theorem. It formalizes key concepts including knowledge representation, semantic information, and the auction-theoretic model of LLM inference. Also, we address some philosophical aspects of the reasoning problem, relevant to our study, its generality, scope and consequences for the AI development.
@@ 4.1 Knowledge and Semantic Information
Let us define knowledge as organized information that LLMs can encode and utilize for reasoning, and aligning with epistemological concepts while maintaining mathematical rigor. In other words, knowledge, much like energy (with low entropy) in physics, is a quantity we can use to perform work of reasoning. Information is a (measure of) potential for uncertainty reduction in a given context. We assume finiteness of knowledge representation to refer directly to the properties of the embedding space (knowledge representation space) in LLMs.
We adopt a pragmatist perspective on knowledge, treating it as organized information patterns accessible to computational systems. That differs from the probabilistic credence perspective, natural in Bayesian learning, treating knowledge as justified true beliefs updated based on evidence. That is a rather bold ontological claim, but as we will see, one that allows us to make progress in understanding what LLMs can operationally achieve.
Now, let us present formal definitions of these ideas.
Definition 1 (Knowledge).
Knowledge is a finite metric space $\mathcal{K}$ with distance function $d_{\mathcal{K}}:\mathcal{K}\times\mathcal{K}\rightarrow\mathbb{R}_{+}$ that collects all facts, concepts, and relationships that can be encoded by a computational system.
@@
It is natural to think of the knowledge space as the space of embeddings with additional measure and mappings. For model $\mathcal{M}$ with parameters trained on data, $\mathcal{K}_{M}\subset\mathcal{K}$ represents the knowledge subset encoded in the model’s parameters. The distance function $d_{\mathcal{K}}$ measures differences between knowledge elements. The elements of the knowledge space represent atomic facts, relational structures, and transformation rules at many scales. A pragmatist view is to interpret elements of knowledge as activation patterns (ideas) encoding anything from individual tokens of texts to complex sentences or compositional structures.
@@
There is an important (ontological) assumption here that calls for our attention. We purposefully assume $\mathcal{K}$ is closed under transformations available in computable function space. At first glance, that assumption may look as limiting the applicability of our conclusions to combinatorial inference mechanisms, merely recombining subsets of knowledge. That intuition is misleading. In the following sections we introduce the notion of emergence operator to describe reasoning process capable of making novel insights in that setting.
Definition 2 (Knowledge Representation).
Knowledge representation is a tuple $(\mathcal{K},d_{\mathcal{K}},\mathcal{W},\mathcal{A},\Phi,\Psi)$ where $\mathcal{K}$ is a metric space of knowledge elements with distance function $d_{\mathcal{K}}$, $\mathcal{W}$ is the space of all possible model parameters, $\mathcal{A}$ is the space of all possible activation patterns, $\Phi:\mathcal{W}\rightarrow\mathcal{P}(\mathcal{K})$ maps parameters to knowledge subsets, and $\Psi:\mathcal{A}\rightarrow\mathcal{P}(\mathcal{K})$ maps activations to accessible knowledge.
For parameterization $W\in\mathcal{W}$, $\Phi(W)=\mathcal{K}_{M}$ represents encoded knowledge. For activation pattern $a_{i}\in\mathcal{A}$ with parameters $W_{i}$, $\Psi(a_{i})\subset\Phi(W_{i})$ represents knowledge accessed during specific inference. Knowledge representation models how LLMs encode information in their parameters and retrieve it from them.
@@
We need to introduce an idea of computational independence that allows for distinguishing between facts derivable from other facts and facts that cannot be derived that way. We need that distinction to deal with the mechanics of reasoning process.
@@
Let
$\mathcal{F}_{C}=\{f:\mathcal{P}(\mathcal{K})\to\mathcal{P}(\mathcal{K})\mid\text{cost}(f)\leq C\}$ (1)
denote the set of all computational transformations with computational cost bounded by $C$. Knowledge sets $A,B\subseteq\mathcal{K}$ are computationally independent with respect to computational budget $C$, denoted $A\perp_{C}B$, if and only if for all $f\in\mathcal{F}_{C}$:
@@
To gain some intuitions about $\mathcal{F}_{C}$, let us consider several examples. One is the set of all derivation rules in first-order logic, like modus ponens or universal instantiation, with at most $C$ inference steps. Then, give an axiom set $A\subset\mathcal{K}$, we can produce new subset of knowledge with at most $C$ applications of the admitted inference rules. For example, given $A=\{P,P\to Q\}$ and $B=\{R,R\to S\}$ and the budget $C=2$ for two inference steps, we can derive $\{P,P\to Q,Q\}$ $\{R,R\to S,S\}$ with any $f\in\mathcal{F}_{2}$. Then $A\perp_{2}B$ because $f(A)\cap B=\emptyset$ and $A\cap f(B)=\emptyset$, there are no shared elements of knowledge accessible in 2-step derivations.
The computational cost (or complexity) measure $\text{cost}(f)$ could therefore be: Kolmogorov complexity giving the shortest program computing $f$, longest path in computation graph (or length of the chain of thoughts CoT), number of inference rules applied, or number of QKV matrix multiplications in transformers. Then we can define a context-relevant subset of knowledge that can contribute to reasoning within a given context (of a query) within budget some computational budget.
@@
The context-relevant knowledge $\mathcal{R}_{C}(Q)$ for context $Q\subseteq\mathcal{K}$ within computational budget $C$ is:
$\mathcal{R}_{C}(Q)=\{k\in\mathcal{K}:\{k\}\not\perp_{C}Q\}.$ (3)
We can now introduce the concept of semantic information measure to describe a general process of reasoning. That special mapping measures information available in any subset of knowledge and extracts its meaning within available computational budget.
@@
$\mu_{C}:\mathcal{P}(\mathcal{K})\times\mathcal{P}(\mathcal{K})\rightarrow\mathbb{R}_{+}$ (4)
parameterized by computational budget $C$ satisfying:
$\mu_{C}(\emptyset|Q)=0\text{ for all }Q\subseteq\mathcal{K}$ (null set)
$\mu_{C}(A|Q)\geq 0\text{ with }\mu_{C}(A|Q)>0\text{ iff }A\cap\mathcal{R}_{C}(Q)\neq\emptyset$ (contextual non-negativity)
$\mu_{C}(A|Q)\leq\mu_{C}(B|Q)\text{ if }A\subseteq B$ (monotonicity)
$\mu_{C}(A\cup B|Q)\leq\mu_{C}(A|Q)+\mu_{C}(B|Q)\text{ with equality for }A\perp_{C,Q}B$ (subadditivity)
$\mu_{C}(A|Q)\leq\mu_{D}(A|Q)\text{ if }C<D$ (insight monotonicity)
$\sup\{\mu_{C}(A|Q):A\subseteq\mathcal{K}\}<\infty$ (boundedness)
where $A\perp_{C,Q}B$ means $A\perp_{C}B$ within $\mathcal{R}_{C}(Q)$.
Knowledge that cannot resolve any uncertainty or knowledge irrelevant to a query provides no information. However, if there is uncertainty resolving information within our cognitive reach, we can use it and perform reasoning organizing that information into available knowledge. That is why we assume information content is additive for independent knowledge sets, i.e., when no knowledge in $A$ can be derived from $B$ within computational budget $C$, and vice versa. With redundant information in $A\cup B$ , we admit potential loss of information that may be caused by aggregation. Finally, with increasing computational budget we can extract more knowledge (ordered information) from the same information subset.
@@
The choice of conditional subadditivity for computationally independent sets, i.e., demanding $\mu_{C}(A\cup B|Q)\leq\mu_{C}(A|Q)+\mu_{C}(B|Q)$ with equality for $A\perp_{C}B$, is necessary to describe LLM inference as an auction of ideas. We need to isolate knowledge encoded in distinct neural circuits, attention heads, or activation patterns. The concept of independence is the formal translation of this informational separation into a computational context. It states that the knowledge in idea $A$ cannot be used to derive the knowledge in idea $B$ within a given computational budget $C$, and vice-versa.
Given this premise of informational isolation, the additivity of the semantic measure becomes the baseline definition of non-interaction. If two knowledge sets are truly independent, they offer no redundant or synergistic information. That baseline is necessary to identify and quantify deviations. As we will see in the following section, the aggregation of diverse probabilistic beliefs violates this additivity, representing a form of forced synergy imposed by the inference process.
However, there is also another form of synergy that we have already included in our model. To see what it may be, we should understand how context $Q$ may affect the semantic measure of any give subset $A$ of knowledge. The semantic information measure can exhibit superadditivity in the context argument.
To see that, suppose $Q_{1}=\{P\}$ and $Q_{2}=\{P\to R\}$, and consider $A=\{P\}$. If computational budget $C=1$ allows one step and we have $f\in\mathcal{F}_{C}$ representing modus ponens, then
$R\cap f(Q_{1})=\emptyset$ $\text{ and }f(R)\cap Q_{1}=\emptyset,$
$R\cap f(Q_{2})=\emptyset$ $\text{ and }f(R)\cap Q_{2}=\emptyset,$
$R\cap f(Q_{1}\cup Q_{2})\neq\emptyset$ $\text{ and }f(R)\cap(Q_{1}\cup Q_{2})\neq\emptyset.$
@@
$\mu_{C}(A|Q_{1}\cup Q_{2})>\mu_{C}(A|Q_{1})+\mu_{C}(A|Q_{2})=0.$
Integrating different pieces of context may synergistically unlock reasoning paths not available individually.
Another important consequence of the assumptions we have made is that we permit structures where $\mu_{C}$ is not monotonic in context. Consider, for the sake of example, $\mu_{C}(A|Q)=|A\setminus Q|$ and $\mathcal{R}_{C}(Q)=\mathcal{K}\setminus Q$. It is rather easy to see that $\mu_{C}$ satisfies the axioms of the semantic information measure. Let us then see what happens if $A=\{k\}$, $Q_{1}=\emptyset$ and $Q_{2}=\{k\}$. We have:
$Q_{1}\subseteq Q_{2}\quad\text{and}\quad\mu_{C}(A|Q_{1})=|\{k\}\setminus\emptyset|=1>0=|\{k\}\setminus\{k\}|=\mu_{C}(A|Q_{2}).$
Increasing the context may reduce the information contributions. If the extended context $(Q_{2})$ provides an alternative explanation for phenomena related to original context $(Q_{1})$, then marginal contribution of available knowledge $(A)$ may decrease.
The axiomatic choices we have made are most applicable to the idealized setting of Theorem 6. Later, in Theorem 8 and 9, we relax them by moving the source of interaction from the measure itself to the aggregation function in actual transformers.
4.1.1 Limits of Knowledge and Reasoning

### v7

@@ 4 Theoretical Framework
This section develops the mathematical foundations necessary for proving the Impossibility Theorem. It formalizes key concepts including knowledge representation, semantic information, and the auction-theoretic model of inference. Also, we address some philosophical aspects of the reasoning problem, relevant to our study, its generality, scope and consequences for the AI development.
@@ 4.1 Knowledge and Semantic Information
If we define knowledge as organized information that AI models can encode and utilize for reasoning, then, much like low-entropy energy in physics, it becomes a quantity used to perform the work of reasoning. Information, in turn, represents the potential to reduce uncertainty in a given context. This pragmatist perspective treats knowledge as a space of organized patterns accessible to computational transformations. This is the perspective we will adopt, and we will now present formal definitions of these ideas.
4.1.1 Knowledge lives in Polish space
To describe knowledge in a general way, one that is well suited for our study of inference processes, we need a structure with properties reflecting the reality we describe. First, for modeling and technical reasons, we need a way to measure distance between any pair of elements of knowledge. That is provided by a complete metric space. It is also reasonable to demand that each element of knowledge has a neighborhood consisting of countably many knowledge elements that are arbitrary close, to make exploration or understanding of ideas possible. Therefore, the space we need must be complete to guarantee the convergence of continuous transformations (inherent in neural networks) or to support the existence of limits required to model the stabilization of reasoning processes.
In other words, we need a complete, separable metric space, i.e., a Polish space kuratowski1966wstep .
Definition 1 (Polish space).
A topological space is called Polish if it is a complete metric space that has a countable dense subset.
That space is sufficiently general to include Banach spaces, Euclidean distance, or continuous mappings, and provides the necessary topological properties for developing advanced (probability) measure theory. It is also relevant to reasoning processes that we are going to describe as computational transformations. Throughout, $(\mathcal{K},d_{\mathcal{K}})$ denotes a Polish (complete, separable metric) space.
Reasoning over a subset of knowledge generates another (possibly new) subset of knowledge. That requires operations in spaces that are closed under set-theoretical transformations, including countable unions and intersections. Borel $\sigma$-algebras of open sets of $\mathcal{K}$, denoted $\mathcal{B}(\mathcal{K})$, support development of such consistent knowledge generation and measurement processes. Then, the necessary closure properties for modeling reasoning are provided by the family $\mathcal{A}(\mathcal{K})$ of analytic subsets of $\mathcal{K}$. Analytic subsets of Polish spaces are continuous images of Borel subsets, closed under countable unions and intersections.
Finally, to define a measure of information provided by any subset of knowledge, we utilize the $\sigma$-algebra of universally measurable sets, denoted $\mathcal{U}(\mathcal{K})$, i.e., the $\sigma$-algebra of sets measurable w.r.t. every Borel probability measure on $\mathcal{K}$.
It follows from the descriptive set theory that $\mathcal{B}(\mathcal{K})\subset\mathcal{A}(\mathcal{K})\subset\mathcal{U}(\mathcal{K})$, see e.g. Kechris1995 . Analytic sets are closed under countable unions, Borel-measurable images, and intersections with Borel sets. Arbitrary intersections of analytic sets need not be analytic in general, but analytic sets are universally measurable. Differences of analytic sets need not be analytic, but they are universally measurable, as well.
These properties ensure that reasoning operations are well-defined and remain within a measurable domain. That is the structure rich enough for our needs. To justify that choice, let us investigate limitations of simpler structures.
Example.
Let $X$ be a connected subset of a Polish space. Consider a continuous mapping $\Phi\colon X\mapsto\mathcal{K}$ modeling inference based on self-attention and MLP.
If $\mathcal{K}$ is finite, and hence discrete, then by continuity the image $\Phi(X)$ is connected in a discrete space. That means it must be constant. As a result, any nontrivial continuous dependence on embeddings is impossible if $\mathcal{K}$ is finite. A Polish subspace removes this problem. ∎
Example.
For any two probability distributions $\pi^{(1)},\pi^{(2)}\in\Delta^{\infty}$ (over countable dictionaries) and $\beta\in(0,1)$, the convex aggregate $\Pi=\beta\pi^{(1)}+(1-\beta)\pi^{(2)}$ defines the line segment $\{\beta\pi^{(1)}+(1-\beta)\pi^{(2)}:\beta\in[0,1]\}$. That is an uncountable subset of the Polish space. Therefore, knowledge space closed under such aggregation cannot be finite. In contrast, it resides naturally in a Polish domain. ∎
We now formalize the knowledge space and the reasoning transformations acting upon it.
Definition 2 (Knowledge Domain and Relevance).
Knowledge is a Polish space $(\mathcal{K},d_{\mathcal{K}})$ in which all reasoning sets are analytic subsets of $\mathcal{A}(\mathcal{K})$.
@@ 4.1.1 Knowledge lives in Polish space
It is natural to think of the knowledge space as the space of embeddings with additional measure and mappings. For model $\mathcal{M}$ with parameters trained on data, $\mathcal{K}_{M}\subset\mathcal{K}$ represents the knowledge subset encoded in the model’s parameters. The distance function $d_{\mathcal{K}}$ measures differences between knowledge elements. The elements of the knowledge space represent atomic facts, relational structures, and transformation rules at many scales. We need a continuous, complete metric knowledge space. Without that requirement, we will not be able to model the internal numerical process by which the model reasons and transforms the input symbols into output symbols.
@@
We make an important (ontological) assumption that calls for attention and explanation. We assume that knowledge is closed under transformations available in computable function space. That suggests we are merely recombining subsets of knowledge. As we will see in the following sections, that intuition is misleading since the reasoning processes are very much capable of making novel insights in that setting.
4.1.2 Reasoning envelope
We model the reasoning process using a set of computational primitives. These create a reasoning envelope and define the scope of our inference studies.
Definition 3 (Reasoning Envelope).
For each computational budget $C>0$, let
$\mathcal{F}_{C}=\{f\colon\mathcal{K}\to\mathcal{K}\mid\text{cost}(f)\leq C\}$ (1)
denote a countable family of Borel measurable computable functions. We assume composition-closure with cost additivity, i.e., if $f\in\mathcal{F}_{C}$ and $g\in\mathcal{F}_{D}$, then $g\circ f\in\mathcal{F}_{C+D}$, and $\mathcal{F}_{C}\subseteq\mathcal{F}_{D}$ whenever $C<D$. The total computability envelope is
$\mathcal{F}_{\infty}=\bigcup_{C\geq 0}\mathcal{F}_{C}.$ (2)
$\mathcal{F}_{C}$ is a list of functions that are well-behaved (Borel measurable), efficient to compute (bounded cost), and if any two of them, say $f$ and $g$, form a composition $h(A)=(f\circ g)(A)=f(g(A))$, then $h$ is also on the list (closed under composition).
Because $f\in\mathcal{F}_{C}$ is Borel measurable, the image $f(A)$ of any analytic knowledge subset $A\in\mathcal{A}(\mathcal{K})$ is also analytic Kechris1995 . This guarantees that the reasoning process maps analytic sets to analytic sets, as desired.
Examples of computational cost (or complexity), $\text{cost}(f)$, include the Kolmogorov complexity kolmogorov1963tables giving the shortest program computing $f$, the longest path in computation graph (or length of the chain of thoughts CoT), a number of inference rules applied, or a number of QKV matrix multiplications in transformers.
Example.
Consider a reasoning primitive calculating an linear combination of any pair of vectors, $f\colon\{v_{1},v_{2}\}\to\{y(t):t\in\mathbb{R}\}$. For $t\in\mathbb{R}$ and attention score $\alpha(t)=1/(1+e^{-t})$, we then have
$y(t)=\alpha(t)v_{1}+\bigl(1-\alpha(t)\bigr)v_{2}.$ (3)
A finite knowledge space $\mathcal{K}$ cannot contain all such $y(t)$ while remaining closed under the transform $f$.
By contrast, taking $\mathcal{K}$ to be a Polish space that contains $y(t)$, e.g., $\mathbb{R}^{d}$ with Euclidean metric, accommodates such attention outputs. ∎
@@ 4.1.2 Reasoning envelope
The following idea of computational independence allows for distinguishing between facts derivable from other facts and facts that cannot be derived that way. We need that distinction to deal with the mechanics of reasoning process.
@@
Knowledge sets $A,B\subseteq\mathcal{K}$ are computationally independent with respect to computational budget $C$, denoted $A\perp_{C}B$, if and only if for all $f\in\mathcal{F}_{C}$:
@@
To gain some intuitions about $\mathcal{F}_{C}$ and independence, consider the set of all derivation rules in first-order logic, like modus ponens or universal instantiation, with at most $C$ inference steps. Then, give an axiom set $A\subset\mathcal{K}$, we can produce new subset of knowledge with at most $C$ applications of the admitted inference rules.
Example.
Given
$A=\{P,P\to Q\}\text{ and }B=\{R,R\to S\},$ (5)
and the budget $C=1$ for one inference step, we can derive
$\{P,P\to Q,Q\}\text{ and }\{R,R\to S,S\}$ (6)
with any $f\in\mathcal{F}_{1}$. Then $A\perp_{1}B$, i.e., sets are independent because $f(A)\cap B=\emptyset$ and $A\cap f(B)=\emptyset$. There are no shared elements of knowledge accessible in 1-step derivations. ∎
Since reasoning is performed within a context, we define context-relevant subset of knowledge that can contribute to reasoning within some computational budget.
@@
The context-relevant knowledge $\mathcal{R}_{C}(Q)$ for context $Q\in\mathcal{B}(\mathcal{K})$ within computational budget $C>0$ is:
$\mathcal{R}_{C}(Q)=\{k\in\mathcal{K}:\{k\}\not\perp_{C}Q\}\in\mathcal{A}(\mathcal{K}).$ (7)
Based on the definition above, we say that knowledge $\{k\}$ is relevant when
[…]
There is something we can derive from that knowledge that has something in common with a given context, or we can find a relation with what we know when we reason about that context. The context-relevant knowledge $\mathcal{R}_{C}(Q)$ is analytic because it is a countable union of Borel sets and the Borel context $Q$.
4.1.3 The Semantic Information Measure
We now introduce the concept of semantic information measure to describe a general process of reasoning. That special mapping measures information available in any subset of knowledge and extracts its meaning within available computational budget.
@@ 4.1.3 The Semantic Information Measure
$\mu_{C}\colon\ \mathcal{U}(\mathcal{K})\times\mathcal{B}(\mathcal{K})\mapsto\mathbb{R}_{+},$ (9)
parameterized by computational budget $C>0$ satisfying the following axioms:
(null set) $\mu_{C}(\emptyset|Q)=0\text{ for all }Q\subseteq\mathcal{K},$ (10)
(contextual non-negativity) $\mu_{C}(A|Q)\geq 0\text{ with }\mu_{C}(A|Q)>0\text{ iff }A\cap\mathcal{R}_{C}(Q)\neq\emptyset,$ (11)
(monotonicity) $\mu_{C}(A|Q)\leq\mu_{C}(B|Q)\text{ if }A\subseteq B,$ (12)
(subadditivity) $\mu_{C}(A\cup B|Q)\leq\mu_{C}(A|Q)+\mu_{C}(B|Q),$ (13)
(insight monotonicity) $\mu_{C}(A|Q)\leq\mu_{D}(A|Q)\text{ if }C<D,$ (14)
(boundedness) $\sup\{\mu_{C}(A|Q):A\in\mathcal{U}(\mathcal{K})\}<\infty.$ (15)
Knowledge that cannot resolve any uncertainty or knowledge that is irrelevant to the given context provides no information. If there is an uncertainty resolving information within our cognitive reach, then we can use it and perform reasoning organizing that information into available knowledge. We assume information content is subadditive for independent knowledge sets. When no knowledge in $A$ can be derived from $B$ within computational budget $C$, and vice versa. With redundant information in $A\cup B$ , we admit potential loss of information that may be caused by aggregation. With increasing computational budget we can extract more knowledge (ordered information) from the same information subset.
In the absence of context, when $Q=\emptyset$, nothing is relevant so $\mu_{C}(A|\emptyset)=0$. Knowledge provides zero semantic information, is useless, when there is no context defining the uncertainty to be resolved. That means, we strongly emphasize the contextual relativity of information. Knowledge is only informative relative to a specific query or problem.
We emphasize that the axioms require $\mu_{C}$ to be subadditive with respect to the knowledge argument $A$. Its behavior with respect to the context $Q$ is distinct and permits superadditivity, modeling how integrating contexts can unlock synergistic insights (see Example below).
The axiom of boundedness reflects the fundamental constraint that bounded reasoning systems can only access bounded information, even within an infinite knowledge space. The axiom does not contradict the potential infinity of the knowledge space itself. It asserts that the accessible information content derived from this space must remain finite. By the monotonicity axiom, this is also equivalent to stating that the measure of the entire knowledge space is finite, $\mu_{C}(\mathcal{K}|Q)<\infty$.
@@
The choice of conditional subadditivity for computationally independent sets is necessary to describe isolated knowledge encoded in distinct neural circuits, attention heads, or activation patterns. The concept of independence is the formal translation of this isolation into a computational context. If two knowledge sets are truly independent, they offer no redundant or synergistic information. The subadditivity axiom becomes the baseline definition of non-interaction.
However, there is also a form of synergy already included in the model. To see what it may be, we should understand how context $Q$ may affect the semantic measure. The semantic information measure can exhibit superadditivity in the context argument. Integrating different pieces of context may unlock reasoning paths not available individually.
Example.
Suppose $Q_{1}=\{P\}$ and $Q_{2}=\{P\to R\}$, and consider $A=\{P\}$. If computational budget $C=1$ allows one step and we have $f\in\mathcal{F}_{C}$ representing modus ponens, then
$R\cap f(Q_{1})=\emptyset$ $\text{ and }f(R)\cap Q_{1}=\emptyset,$ (16)
$R\cap f(Q_{2})=\emptyset$ $\text{ and }f(R)\cap Q_{2}=\emptyset,$ (17)
$R\cap f(Q_{1}\cup Q_{2})\neq\emptyset$ $\text{ and }f(R)\cap(Q_{1}\cup Q_{2})\neq\emptyset.$ (18)
@@
$\mu_{C}(A|Q_{1}\cup Q_{2})>\mu_{C}(A|Q_{1})+\mu_{C}(A|Q_{2})=0.$ (19)
Those axiomatic choices are most applicable to the idealized setting of Theorem 6. Later, in Theorem 8 and 9, we relax them by moving the source of interaction from the measure itself to the aggregation function in actual transformers.
4.1.4 Limits of Knowledge and Emergence Operator